# Augmentation-based Unsupervised Cross-Domain Functional MRI Adaptation for Major Depressive Disorder Identification

### v1

@@ IV-B2 Domain Adaptation Module
To reduce domain offset, we employ a maximum mean difference constraint to align the learned feature on different domains. As shown in Fig. 1, we obtain the graph-level representation $f_{X^{d}}$ of the source and target domains through the Transformer module learning. The $f_{X^{d}}$ are then fed into two fully connected layers, which are finally predicted by a softmax layer.
Here, considering that the low-level features contain limited semantic information, we only align the graph-level feature mappings on the source and target domains, represented as follows:
$L_{M}=\frac{1}{B^{2}}\sum_{l=1}^{L}\sum_{i=1}^{n}\sum_{j=1}^{n}(f_{s}^{l(i)}-f_{t}^{l(j)})^{T}(f_{s}^{l(i)}-f_{t}^{l(j)}),$ (5)
where $f_{s}^{l(i)}$ is the features of the $i$-th sample in the source domains, $f_{t}^{l(j)}$ are the features of the $j$-th sample in the target domains, $B$ is the batch size, $n$ is the number of samples, and $L$ denotes the number of layers in the fully connected layer. Eq. (5) measures the discrepancy in feature distribution between the two domains. By minimizing this equation, we penalize the features with different distributions across domains, compelling the Transformer feature extractor to learn shareable feature representations across the two domains.
@@ IV-B3 Augmentation-based Self-optimization Module
In Section IV-B1, we utilize a Transformer encoder to learn the graph representation for each subject. However, due to the limited training data and the large number of parameters in the multilayered self-attention mechanism and the feed-forward neural network, the model is prone to overfitting. In the unsupervised domain adaptation setting, the risk of the model overfitting the source domain data is significantly heightened as a result of the model’s reliance solely on supervision from the source domain labels. Therefore, we reduce the risk of model overfitting on the source domain data by augmenting the target domain data to improve model generalizability.
To enhance model robustness, we randomly select the input features of the $l$-th layer Transformer encoder for augmentation in the target domain. Given the initial graph feature $X^{t}$ to be augmented, we utilize another randomly selected subject $X^{t^{\prime}}$ from the target domain to add an offset to $X^{t}$ in order to achieve augmentation. Let $Z^{l}_{X^{t}}$, $Z^{l}_{X^{t^{\prime}}}$ denote the input features of $X^{t}$ and $X^{t^{\prime}}$ in layer $l$, respectively. The features of subject $X^{a}$ in the $l$-th layer after augmentation can be expressed as:
$Z^{l}_{X^{a}}=Z^{l}_{X^{t}}+\alpha(Z^{l}_{X^{t^{\prime}}}-{Z}^{l}_{X^{t}}),$ (6)
where $\alpha$ is a scalar that is used to adjust the degree of offset to the target domain features $X^{t}$. In fact, such an augmentation operation is equivalent to imposing regularization constraints on the model. It is worth noting that by randomly selecting the $l$-th layer for augmentation, the learning of subsequent parameters from the $(l+1)$-th layer to the last layer will be affected, thus achieving a multilayer regularization constraint on the model. As shown in Fig. 1, $Z^{l}_{X^{t}}$ is augmented to the new target domain data $Z^{l}_{X^{a}}$ after the perturbation of $Z^{l}_{X^{t^{\prime}}}$ in the $l$-th layer. Then, the new feature representations $Z_{X^{t}}$ and $Z_{X^{a}}$ are obtained through the subsequent learning of the Transformer layer. We calculate the prediction probabilities of $Z_{X^{t}}$ and $Z_{X^{a}}$ by performing the same procedure as mentioned (i.e., flatten, fully connected layers and softmax), and denoted as $P_{X^{t}}$, $P_{X^{a}}$. Then, we compute the Kullback-Leibler ($KL$) divergence between $P_{X^{t}}$ and $P_{X^{a}}$ to minimize it:
@@
It is worth noting that we do not have label information for the target domain to determine whether predictions about the target domain are reliable. Therefore, a confidence filter is added to round off samples corresponding to prediction probabilities with low confidence. The set of remaining target domain data can be represented as:
@@ IV-B4 Classification
The resulting graph-level representation of each subject is fed into two fully connected layers for classification, and the prediction probability vector is obtained through the softmax layer. The number of neurons in the first and second layers was $4096$ and $2$, respectively. The whole model is trained using a joint loss function $L$ as the optimization objective, which comprises three components: (1) cross-entropy loss $L_{C}$ measures the difference between the model’s prediction for source domain data and the real label to ensure the model’s classification performance in source domain, (2) maximum average difference loss $L_{M}$ is used to align the learned inter-domain features, (3) augmentation-based self-optimization loss $L_{A}$ is employed to further optimize the model and prevent overfitting of the model on the source domain data. Therefore, the joint loss function can be expressed as:
@@ IV-C Implementation Details
The proposed AUFA model is implemented on Pytorch using NVIDIA Quadro RTX 6000. The model is trained in two stages. In the first stage, the AUFA is trained in a supervised manner based on the source domain to narrow down the parameter search space. In the second stage, the pre-trained AUFA is further optimized by minimizing the joint loss function Eq. (10) based on all labeled source domain data and unlabeled target domain data. We set the number of layers in the Transformer encoder to $2$ and the number of attention heads to $4$. Hyperparameters $\lambda_{1}$ and $\lambda_{2}$ are both $1$. The AUFA model is optimized by Adam algorithm with a learning rate of $0.00001$, training epoch of $45$ and batch size of $32$.
TABLE II: Results (mean±$\pm$±standard deviation) of eight methods in MDD vs. NC classification, and the best results are shown in bold.
@@
V Experiments
V-A Experimental Settings
In this work, we designed 4 cross-site fMRI prediction tasks: SITE 20 $\rightarrow$ SITE 21, SITE 20 $\rightarrow$ SITE 25, SITE 20 $\rightarrow$ SITE 1, SITE 20 $\rightarrow$ SITE 9. The dataset in front of the arrow is defined as the source domain, and the dataset after the arrow is defined as the target domain. The source domain dataset contains full category labels and the target domain dataset does not have any labeling information. Considering the limited number of samples, we use all source/target domain samples to train and test the categories to which all target domain samples belong. We trained the model with 5 random initializations and recorded the mean and standard deviation results. The model was evaluated using 5 metrics, including (1) accuracy, (2) precision, (3) recall, (4) area under the receiver operating characteristic curve (AUC), and (5) F1-score.
@@ V-B Competition Methods
We compare the proposed AUFA with three traditional (BC-SVM, LE-SVM and MNSFS) and four deep learning methods (GIN, ST-GCN, CORAL, and DANN). Among them, BC-SVM, LE-SVM, MNSFS, GIN and ST-GCN have not undergone adaptive learning, meaning that the model trained on the source domain data is directly used to predict the target domain data. Note that CORAL, DANN and AUFA are different domain adaptation methods, and unlabeled target domain data will be used for training a specific model.
@@
(1) BC-SVM: In this method, we construct an FCN for each subject based on PC. The betweenness centrality (BC) [44] of each node (ROI) is then extracted from the FCN to measure the number of shortest paths between that node and other nodes in the network. Finally, the corresponding FCN for each subject has a 116-dimensional feature and is classified by the SVM.
@@
(2) LE-SVM: Similar to BC-SVM, we extract the local efficiency [45] of each node as features ($116$ dimension) based on the constructed FCN, followed by SVM for classification.
@@
(3) MNSFS: The method extracts multiple measures for each node from the estimated FCN, including three different definitions of local clustering coefficients, four centralities (i.e., degree centrality, betweenness centrality, closeness centrality, and eigenvector centrality), as well as local efficiency. The measures belonging to each node are assigned to a group, followed by sparse group LASSO for feature selection and SVM for classification [46].
@@
(4) GIN: It uses graph isomorphic network (GIN) [47] for FCN feature extraction. Each GIN layer aggregates the neighbor features of each node based on FCN and uses MLP to perform nonlinear transformation to achieve node update. This process can be repeated through multiple layers of the same structure to gradually fuse the global information. Finally, the output of the GIN module is average pooled to obtain the graph representation features for each subject, and classify them by the fully connected layer. The number of GIN layers is $2$, and the output dimension of each layer is $64$.
@@
(5) ST-GCN: It models dynamic features of fMRI through spatio-temporal graph convolutional networks (ST-GCN) [48]. We input regional fMRI time series into two spatio-temporal graph convolutional units to learn the spatio-temporal representation of fMRI. Finally, the classification tasks of MDD and NC are performed through fully connected layers. The output dimension of each convolutional unit is $64$.
@@
(6) CORAL: In this method, we use the same graph representation learning method as the proposed AUFA to extract features from each subject’s FCN. Here, we use the label of the source domain for supervised training, and apply the covariance alignment (CORAL)  [49] module to minimize the distribution differences between the source and target domain data, thereby reducing the effect of domain offset.
@@
(7) DANN: The domain adversarial neural network (DANN) [50] is a deep learning model for domain adaptation. In DANN, a convolutional neural network is used to extract the features of the source and target domains, and the domain discriminator is used to judge which domain the data comes from. Similar to AUFA, DANN only has supervision of source domain and learns domain invariant features through adversarial training of domain discriminators and classifiers.
@@ V-C Experimental Results
In Table II, we report the experimental results of our proposed AUFA and $7$ comparison methods in MDD identification tasks. Statistical significance analysis results are reported in Supplementary Materials. From the Table II, we can make the following observations. First, the methods employing domain adaptive strategies (i.e., CORAL, DANN, and our proposed AUFA) exhibit significant advantages when compared to learning methods without domain adaptive processing (i.e., BC-SVM, LE-SVM, MNSFS, GIN, and ST-GCN). This phenomenon reveals the limitation on the generalization performance of models trained directly from the source domain to the target domain due to the significant data heterogeneity problem among different sites. This validates the effectiveness of utilizing the maximum mean difference constraint to achieve inter-domain feature alignment. Second, deep learning models such as GIN and ST-GCN show relatively high performance in methods that do not perform domain adaptation. This is likely due to their capability to adaptively learn discriminative features that are beneficial for the classification task under label supervision. In contrast, traditional machine learning methods heavily rely on predefined features and prior knowledge. Furthermore, our AUFA model achieves excellent performance on most of the metrics in several domain adaptive methods. Compared with CORAL, which adopts the same graph representation learning in our method, AUFA achieves an average improvement of about 5% and 4% in ACC and AUC metrics on the four domain adaptation tasks. This improvement can be attributed to the augmentation-based self-optimization module integrated within AUFA, which effectively alleviates the problem of model overfitting on the source domain, resulting in beneficial optimization of the overall model performance. Finally, compared to the popular DANN method, AUFA still maintains outstanding performance with an average improvement of about 5.5% and 5% in ACC and AUC metrics across the four domain adaptation tasks. Besides the augmentation of target data by AUFA, which mitigates the risk of overfitting, another possible reason is that the convolutional neural network structure adopted by DANN may not fully recognize and focus on the importance of features from different brain regions. In contrast, AUFA leverages self-attention mechanisms to learn different weights for each brain region, thus extracting more discriminative features.
@@ V-D Visualization of FMRI Data Distribution
We visualize the distribution differences between the features at each site before and after domain adaptation. Specifically, we flattened the upper triangular matrix of the FCNs of the source and target domain data as the features without domain adaptation, and then, the features learned by the Transformer in our AUFA framework as the features after domain adaptation. Finally, these features are embedded into a 2D space for visualization by t-SNE [43], with results displayed in Fig. 3. As can be seen in Fig. 3, the feature distributions of the two sites that have not undergone domain adaptation show a large domain offset. After domain adaptation, the distributions of the source and target domains become consistent. This indicates that our method can effectively reduce the heterogeneity of the fMRI data from different sites and thus improve the performance of the model on the target domain.
@@ VI-A Ablation Experiment
To explore the effectiveness of the AUFA framework more deeply, we designed three experiments to further analyze the effect of each sub-module: (1) Feature extraction is performed using the same backbone network as AUFA, but without any domain adaptation and augmentation operations (only $L_{C}$ is retained), i.e., AUFA-C; (2) The domain adaptation module in the AUFA framework is discarded and only $L_{C}$ and $L_{A}$ are retained, i.e., AUFA-AUG; (3) Without augmenting the data in the target domain, only $L_{C}$ and $L_{M}$ are retained, i.e., AUFA-MMD. We conduct ablation experiments on the largest sample classification task (i.e., SITE 20 $\rightarrow$ SITE 21), and the results are shown in Table III. It can be observed that our AUFA obtains the best classification results, and removing any part of the framework results in a degradation of performance. This demonstrates the effectiveness of our proposed framework, in which the maximum mean difference constraint effectively achieves inter-domain feature alignment. Additionally, the augmentation of the target domain alleviates overfitting and improves model generalizability.
TABLE III: Classification results of our AUFA and its three variants in the MDD and NC classification task for cross-domain SITE 20 $\rightarrow$ SITE 21, with the best results shown in bold.

### v2

@@ IV-B2 Domain Adaptation Module
To reduce domain offset, a maximum mean difference constraint is utilized to align the learned features across different domains. As shown in Fig. 1, we learn the graph-level feature $f_{X^{d}}$ of the source and target domains through the Transformer module learning. The $f_{X^{d}}$ are then fed into two fully connected layers, which are finally predicted by a softmax layer.
Here, considering that the low-level features contain limited semantic information, we only align the graph-level feature mappings between domains, represented as follows:
$L_{M}=\frac{1}{B^{2}}\sum_{l=1}^{L}(f_{s}^{l}-f_{t}^{l})^{T}(f_{s}^{l}-f_{t}^{l}),$ (5)
where $f_{s}^{l}$ represents the feature of the $l$-th layer in the source domain, $f_{t}^{l}$ represents the feature of the $l$-th layer in the target domain, $B$ is the batch size, and $L$ denotes the number of layers in the fully connected layer. Eq. (5) measures the discrepancy in feature distribution between the two domains. By minimizing this equation, we penalize the features with different distributions across domains, compelling the Transformer feature extractor to learn shareable feature representations across the two domains.
@@ IV-B3 Augmentation-based Self-optimization Module
In Section IV-B1, we utilize a Transformer encoder to learn the graph representation for each subject. However, the model is prone to overfitting due to the limited training data and the large number of parameters included in the multilayer self-attention mechanism and feed-forward neural network. In the unsupervised domain adaptation setting, the risk of the model overfitting the source domain data is significantly heightened as a result of the model’s reliance solely on supervision from the source domain labels. Therefore, we reduce the risk of model overfitting on the source domain data by augmenting the target domain data to improve model generalizability.
To enhance model robustness, we randomly select the input features of the $l$-th layer Transformer encoder for augmentation in the target domain. Given the initial graph feature $X^{t}$ to be augmented, we utilize another randomly selected subject $X^{t^{\prime}}$ from the target domain to add an offset to $X^{t}$ to achieve augmentation. Let $Z^{l}_{X^{t}}$, $Z^{l}_{X^{t^{\prime}}}$ denote the input features of $X^{t}$ and $X^{t^{\prime}}$ in layer $l$, respectively. The features of subject $X^{a}$ in the $l$-th layer after augmentation can be expressed as:
$Z^{l}_{X^{a}}=Z^{l}_{X^{t}}+\gamma(Z^{l}_{X^{t^{\prime}}}-{Z}^{l}_{X^{t}}),$ (6)
where $\gamma$ is a scalar that is used to adjust the degree of offset to the target domain features $X^{t}$. In fact, such an augmentation operation is equivalent to imposing regularization constraints on the model. Note that by randomly selecting the $l$-th layer for augmentation, the learning of subsequent parameters from the $(l+1)$-th layer to the last layer will be affected, thus achieving a multilayer regularization constraint on the model. As shown in Fig. 1, $Z^{l}_{X^{t}}$ is augmented to the new target domain data $Z^{l}_{X^{a}}$ after the perturbation of $Z^{l}_{X^{t^{\prime}}}$ in the $l$-th layer. Then, the new feature representations $Z_{X^{t}}$ and $Z_{X^{a}}$ are obtained through the subsequent learning of the Transformer layer. We calculate the prediction probabilities of $Z_{X^{t}}$ and $Z_{X^{a}}$ by performing the same procedure as mentioned (i.e., flatten, fully connected layers and softmax), and denoted as $P_{X^{t}}$, $P_{X^{a}}$. Then, we compute the Kullback-Leibler ($KL$) divergence between $P_{X^{t}}$ and $P_{X^{a}}$ to minimize it:
@@
It is important to mention that we lack label information to determine the reliability of predictions about the target domain. Therefore, a confidence filter is added to round off samples corresponding to prediction probabilities with low confidence. The set of remaining target domain data can be represented as:
@@ IV-B4 Classification
The graph-level representation of each subject is fed into two fully connected layers for classification, and the prediction probability vector is obtained through a softmax layer. The number of neurons in the first and second layers was $4096$ and $2$, respectively. The whole model is trained using a joint loss function as the optimization objective, which comprises three components: (1) cross-entropy loss $L_{C}$ measures the discrepancy between the model’s prediction for source domain data and the real label to ensure the model’s classification performance in source domain, (2) maximum average difference loss $L_{M}$ is used to align the learned inter-domain features, (3) augmentation-based self-optimization loss $L_{A}$ is utilized to further optimize the model and mitigate overfitting of the model on the source domain data. Therefore, the joint loss function can be represented as:
@@ IV-C Implementation Details
The AUFA framework is implemented on Pytorch using NVIDIA Quadro RTX 6000. The model is trained in two stages. During the first stage, AUFA is trained in a supervised manner based on the source domain to narrow down the search space for parameters. The pre-trained AUFA is further optimized in the second stage by minimizing the joint loss function Eq. (10) based on labeled source samples and unlabeled target samples. We set the number of layers in the Transformer encoder to $2$ and the number of attention heads to $4$. Hyperparameters $\lambda_{1}$ and $\lambda_{2}$ are both $1$. We optimize the AUFA model using the Adam algorithm with a learning rate of $0.00001$, a training epoch of $45$, and a batch size of $32$.
V Experiments
V-A Experimental Settings
In this work, we design 4 cross-site fMRI prediction tasks: SITE 20 $\rightarrow$ SITE 21, SITE 20 $\rightarrow$ SITE 25, SITE 20 $\rightarrow$ SITE 1, SITE 20 $\rightarrow$ SITE 9. The site in front of the arrow is the source domain and behind the arrow is the target domain. The source domain data contains full category labels and the target domain data does not have any label information. We employ source and target domain samples for training and predict the category to which the target domain samples belong based on the trained model parameters. The model was trained with 5 random initializations, and the mean and standard deviation results were recorded. The model was evaluated using 5 metrics, including (1) accuracy, (2) precision, (3) recall, (4) area under the receiver operating characteristic curve (AUC), and (5) F1-score.
TABLE II: Experimental results (mean±$\pm$±standard deviation) of the eight methods in the MDD vs. NC classification task, with the best results highlighted in bold.
@@ V-B Competition Methods
We compare the proposed AUFA with three traditional (BC-SVM, LE-SVM and MNSFS) and four deep learning methods (GIN, ST-GCN, CORAL, and DANN). Among them, BC-SVM, LE-SVM, MNSFS, GIN and ST-GCN have not undergone adaptive learning, meaning that the model trained on the source domain is directly used to predict category labels of target data. Note that CORAL, DANN and AUFA are different domain adaptation methods, and unlabeled target domain data will be involved in the training of the model.
@@
(1) BC-SVM: In this method, an FCN is constructed for each subject based on PC. We then extract the betweenness centrality (BC) [44] of each ROI (node) from the FCN to measure the number of shortest paths between that ROI and other ROIs in the brain. Finally, the corresponding FCN for each subject has a 116-dimensional feature and is classified by the SVM.
@@
(2) LE-SVM: Similar to BC-SVM, we extract the local efficiency [45] of each node as features ($116$ dimension) based on the constructed FCN, which is then utilized for classification through SVM.
@@
(3) MNSFS: The method extracts multiple measures for each node from the estimated FCN, including local efficiency, three different definitions of local clustering coefficients, and four centralities. The measures belonging to each node are assigned to a group, followed by sparse group LASSO for feature selection and SVM for classification [46].
@@
(4) GIN: It uses graph isomorphic network (GIN) [47] for FCN feature extraction. Each GIN layer aggregates the neighbor features of each node based on FCN and uses MLP to perform nonlinear transformation to achieve node update. This process can be repeated through multiple layers of the same structure to gradually fuse the global information. Finally, the output of the GIN module is average pooled to obtain the graph representation features for each subject, and classify them by the fully connected layer. The method includes $2$ layers of GINs, with each layer having an output dimension of $64$.
@@
(5) ST-GCN: It models dynamic features of fMRI through spatio-temporal graph convolutional networks  [48]. We input regional fMRI time series into two spatio-temporal graph convolutional (ST-GC) units to learn the spatio-temporal representation of fMRI. Finally, the classification tasks of MDD and NC are performed through fully connected layers. Each ST-GC unit outputs features with a dimension of $64$.
@@
(6) CORAL: In this method, we use the same graph representation learning method as the proposed AUFA to extract features from each subject’s FCN. Here, we use the labels of the source domains to perform supervised training and apply the covariance alignment (CORAL) [49] module to align the distributional differences between the domains, thus reducing the effect of domain offset.
@@
(7) DANN: The domain adversarial neural network (DANN) [50] is a deep learning model for domain adaptation. In DANN, a convolutional neural network is used to learn representations of the source/target domain, and the domain discriminator is used to judge which domain the representation comes from. Similar to AUFA, DANN only has supervision of source domain and learns domain invariant features through adversarial training of domain discriminators and classifiers.
@@ V-C Experimental Results
The experimental results of our proposed AUFA and $7$ comparison methods in MDD identification tasks are reported in Table II. Statistical significance analysis results are reported in Supplementary Materials. We can make the following observations from the Table II. First, the methods employing domain adaptive strategies (i.e., CORAL, DANN, and our proposed AUFA) exhibit significant advantages when compared to learning methods without domain adaptive processing (i.e., BC-SVM, LE-SVM, MNSFS, GIN, and ST-GCN). This phenomenon reveals the limitation on the generalization performance of models trained directly from the source domain and tested directly in the target domain. It is attributed to the significant data heterogeneity issue across different sites. This validates the effectiveness of utilizing the maximum mean difference constraint to achieve inter-domain feature alignment. Second, deep learning models such as GIN and ST-GCN show relatively high performance in methods that do not perform domain adaptation. This is likely due to their capability to adaptively learn discriminative features that are advantageous for label-supervised classification tasks. In contrast, traditional machine learning methods heavily rely on predefined features and prior knowledge. Furthermore, our AUFA model achieves excellent performance on most of the metrics in several domain adaptive methods. Compared with CORAL, which adopts the same graph representation learning in our method, AUFA achieves an average improvement of about 5% and 4% in ACC and AUC metrics on the four domain adaptation tasks. This improvement can be attributed to the augmentation-based self-optimization module integrated within AUFA, which helps alleviate the problem of model overfitting on the source domain. Finally, compared to the popular DANN method, AUFA still maintains outstanding performance with an average improvement of about 5.5% and 5% in ACC and AUC metrics across the four domain adaptation tasks. Besides the augmentation of target data by AUFA, which mitigates the risk of overfitting, another possible reason is that the convolutional neural network structure adopted by DANN may not fully recognize and focus on the importance of features from different brain regions. In contrast, AUFA leverages self-attention mechanisms to learn different weights for each brain region, thus extracting more discriminative features.
@@ V-D Visualization of FMRI Data Distribution
We visualize the distribution differences between the features at each site before and after domain adaptation. Specifically, we flattened the upper triangular matrix of the FCNs of the source and target domain data as the features without domain adaptation, and then, the features learned by the Transformer in our AUFA framework as the features after domain adaptation. Finally, these features are embedded into a 2D space for visualization by t-SNE [43], with results displayed in Fig. 3. As shown in Fig. 3, the feature distributions of the two sites that have not undergone domain adaptation show a large domain offset. After domain adaptation, the distributions of corresponding sites in the source and target domains become consistent. This suggests that our AUFA can effectively mitigate the heterogeneity in fMRI data from different sites, thereby enhancing the model’s performance on the target domain.
@@ VI-A Ablation Experiment
To explore the effectiveness of the AUFA framework more deeply, we design three experiments to further analyze the effect of each sub-module: (1) Feature extraction is performed using the same backbone network as AUFA, but without any domain adaptation and augmentation operations (only $L_{C}$ is retained), i.e., AUFA-C; (2) The domain adaptation module in the AUFA framework is discarded and only $L_{C}$ and $L_{A}$ are retained, i.e., AUFA-AUG; (3) Without augmenting the data in the target domain, only $L_{C}$ and $L_{M}$ are retained, i.e., AUFA-MMD. We conduct ablation experiments on the largest sample classification task (i.e., SITE 20 $\rightarrow$ SITE 21), and the results are reported in Table III. It can be observed that our AUFA obtains the best classification results, and removing any part of the framework results in a degradation of performance. This demonstrates the effectiveness of the AUFA framework, wherein the maximum mean difference constraint effectively achieves inter-domain feature alignment. Additionally, the augmentation of the target domain alleviates overfitting and improves model generalizability.
VI-B Impact of Hyperparameters
[…]
TABLE III: Experimental results of AUFA and its three variants in the MDD and NC classification task for cross-domain SITE 20 $\rightarrow$ SITE 21, with the best results highlighted in bold.